15a.The magnitude of vector A is 29.5 units and points in the direction 330° counterclockwise from the positive x-axis. Calculate the x- and y-components of this vector.


1
Expert's answer
2022-06-06T08:54:33-0400
"A_x=A\cos\theta=29.5\cos330^\circ=25.5"

"A_y=A\sin\theta=29.5\sin330^\circ=-14.8"